OpenStudy (anonymous):

simplify the radical. radical 36a^4b^10

OpenStudy (saifoo.khan):

\[\large 6a^2 b^5\]

OpenStudy (anonymous):

what about radical 300x^16

sam (.sam.):

\[\huge 36a^4b^{10}=(6a ^{2}b ^{5})^{2}=6a ^{2}b ^{5}\]

OpenStudy (saifoo.khan):

\[\large 10x^8 \sqrt3\]

OpenStudy (anonymous):

radical64/81

OpenStudy (saifoo.khan):

\[\sqrt{\frac{64}{81}}\]?

OpenStudy (anonymous):

how do i simplify that??

OpenStudy (saifoo.khan):

cancel out 64 and 81.